# Relaxation time approximation revisited and pole/cut structure in retarded correlators

### v2

@@ III.4 Hard interaction VS soft interaction
$\partial_{t}\chi(x,p)+\bm{v}\cdot\nabla\chi(x,p)=-\mathcal{L}_{0}[\chi]\equiv-v(p)\chi(x,p)+K[\chi],$ (24)
@@
$K[\chi]$ $\equiv f^{1/2}_{eq}(p)p_{0}^{-1}\int{\rm dP^{\prime}}{\rm dP_{1}}{\rm dP_{2}}f_{eq}(p^{\prime})W_{p,p^{\prime}\to p_{1},p_{2}}\big{(}\frac{\chi(x,p_{1})}{f^{1/2}_{eq}(p_{1})}+\frac{\chi(x,p_{2})}{f^{1/2}_{eq}(p_{2})}-\frac{\chi(x,p^{\prime})}{f^{1/2}_{eq}(p^{\prime})}\big{)},$ (26)
where $\nu(p)$ represents the collision frequency and we introduce a compact operator $K$. Here we maintain the same notations as in the previous context to avoid complicating the overall notation system, although the expansion Eq.(23) looks quite different. Furthermore, it is immediately observable that Eq.(24) shares the same form as Eq.(7). As a result, they should also share the same formal solution $\chi(t)=e^{-\mathcal{L}_{0}t}\chi(0)$ when the system is homogeneously perturbed. In both cases, whether considering Eq.(24) or Eq.(7), $\chi$ should be regarded as the perturbation above the same equilibrium state. Given that both Eqs.(24) and (7) are derived from the same Boltzmann equation through linearization, we would not expect the eigenspectrum for $\mathcal{L}_{0}$ in Eq.(24) to differ from that in Eq. (7). If this were the case, the decay behavior of the perturbation would exhibit significant differences, signifying the physical inconsistency. This is another successful application of the observer argument from the previous subsection. Therefore, we argue that the eigenspectrum of $\mathcal{L}_{0}$ should be identical in Eqs.(24) and (7), and this is why we maintain the same notations. Then let’s come to the core of this subsection:
@@
[…]
@@
  • Assume that $\exists 0<\alpha<4,\,\gamma>-2$ and $B^{\prime}>0$, so that $\sigma(g,\Theta)<B^{\prime}g^{-\alpha}\sin^{\gamma}\Theta$, then $\nu(p)<\nu_{0}(p_{0}/m)^{-\epsilon/2}\leq v_{0}$, the interaction is soft, where
@@
    and $\nu_{0}$ is a constant.
Proofs of these theorems can be found in [32]. Analyzing the collision frequency $\nu(p)$ enables us to distinguish between relativistic soft and hard interactions. Mathematical analysis indicates that relativistic interactions tend to be softer than their nonrelativistic counterparts: there is a large class of functions $\sigma(g,\Theta)$, which corresponds to scattering cross-sections for relativistic soft, but non-relativistic hard interactions [31, 32]. It can be also proved that for soft interactions $-\mathcal{L}_{0}$ is a bounded operator with the eigenspectrum $[-\nu_{max},0]$, whereas for hard interactions, $-\mathcal{L}_{0}$ is unbounded, featuring an eigenspectrum of $[-\infty,-\nu_{min}]$, where $\nu_{max}$ and $\nu_{min}$ can be estimated by the extrema of $\nu(p)$. Given the eigenspectrum properties presented here, we conclude that the AW RTA is only well-justified in the case of hard interactions. As most relativistic interactions are soft according to Theorem 2, the AW RTA lacks a solid foundation in the majority of cases. For instance, in the extensively studied scalar $\phi^{4}$ theory, where $\sigma(g,\Theta)\sim\frac{1}{s}=\frac{1}{g^{2}+4m^{2}}<\frac{2}{g^{2}}$, the interaction falls within the soft interaction category (this is confirmed in numerical calculation [13]). This softness results in a branch-cut structure that extends across the entire negative imaginary axis in the retarded correlators [12, 33]. In the next section, the relationship between Q1 and Q2 will be clarified based upon the insights gained from our previous discussions.
III.5 Looking for RTA and eigenspectrum properties of $\mathcal{L}_{\alpha}$
@@
In this subsection, we summarize the approach to identify scenarios where the RTA model is mathematically well-justified. The technical details can be presented as follows. Here, we specify the linearized collision operator in use as $\mathcal{L}_{0}$. According to our discussions about the softness of interactions, the RTA models that are permissible are those based on relativistic hard interactions, although it’s rare for relativistic interactions to be hard. Therefore, looking for RTA turns into testing various systems with different interactions such as scalar $\phi^{4}$ theory, Yukawa theory, QED, QCD etc. For the interaction with simple differential cross section $\sigma(g,\Theta)$, e.g, scalar $\phi^{4}$ theory, we can easily determine its softness with recourse to the mathematical aspects of the linearized Boltzmann equation [31, 32]. For example, in the Weinberg-Salam theory at low energies [23], e.g., four fermions interaction in the electroweak sector well below the gauge boson masses, the typical differential cross section behaves as $\sigma(g,\Theta)\sim s=g^{2}+4m^{2}>\frac{g^{3}}{c_{0}+g}\geq\frac{g^{3}}{c_{0}+g}\sin\Theta$ with the conditions $\gamma=1,\beta=2,B=1$ in Theorem 1. Consequently, the model of a gas consisting of elastically colliding neutrinos in the low-energy limit permits an RTA description of transport phenomena. Whereas for interactions too complex for analytical treatment , or when their classification as hard or soft is not known, we recommend employing the numerical method described in [13] for analyzing the eigenspectrum properties. By implementing the aforementioned procedure, we can establish the dictionary between various field theories and the eigenspectrum structure within their linearized kinetic description so that we can exhaust all commonly used interactions admitting an RTA description.
In III.3, we show that the modified linearized collision operator $\mathcal{L}_{\alpha<0}$ also admits an RTA description provided its nonzero eigenvalues are separated from zero. In this case, the relaxation time takes the form of Eq.(8) but with $E_{p}$ replaced by $m$. However, a physical argument reveals that this class of RTAs is essentially equivalent to the traditional RTA derived from $\mathcal{L}_{0}$. Hence, we choose to neglect this class of RTAs.
The observer argument can not only rule out redundant RTAs, but also constrain certain eigenspectrum properties of $\mathcal{L}_{\alpha\neq 0}$, given the softness of the interactions, as will be detailed below. One may wonder if it is necessary to pay attention to the properties of $\mathcal{L}_{\alpha\neq 0}$ eigenspectrum, especially given that $\mathcal{L}_{0}$ is more naturally associated with the analytical properties of retarded correlators, which will be exhibited in the next section. In principle, the selection of $\mathcal{L}_{\alpha}$ is not fixed; one can select any suitable option. It is more suitable to start from $\mathcal{L}_{\alpha\neq 0}$ sometimes. For instance, a recent calculation on the eigenspectrum of $\mathcal{L}_{1}$ within the massless scalar $\phi^{4}$ theory was given in an analytical way [18], while it is more challenging to perform such calculations for $\mathcal{L}_{0}$ (this may also be the case for other interactions). Therefore, we think that discussing the properties of $\mathcal{L}_{\alpha\neq 0}$ eigenspectrum is not trivial at all.
Then let’s elaborate on how the constraints can be imposed. We opt to present the results via a case-by-case discussion, and remind that “the eigenspectrum ($\gamma_{n}$ or $\gamma^{\prime}_{n}$ for eigenvalues) is gapped or not” and “the decay/relaxation rates, the inverse of the perturbation’s relaxation timescales $\frac{\gamma_{n}}{E_{p}^{\alpha}}$, are gapped or not” are entirely distinct concepts.
  • Soft interactions. The eigenspectrum of $\mathcal{L}_{0}$ extends continuously to the origin, indicating a gapless spectrum.
    • $m=0$: For $m=0$, the relaxation time bound $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ for $\mathcal{L}_{\alpha<0}$ blows up, which is identical to the scenario of $\mathcal{L}_{\alpha>0}$. The relaxation rates are rendered gapless by $E_{p}^{-\alpha}$ , regardless of the eigenspectrum properties or the nature of the interactions. In this case, the observer argument is unconstrained.
    • $m\neq 0$: The eigenspectrum of $\mathcal{L}_{\alpha>0}$ is still gapless for the same reason. For $\alpha<0$, the relaxation timescales are bounded by $\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ provided that the eigenspectrum of $\mathcal{L}_{\alpha<0}$ is gapped. However, to match the gapless nature of $\mathcal{L}_{0}$, the eigenspectrum of $\mathcal{L}_{\alpha<0}$ must be also gapless.
  • Hard interactions. The eigenspectrum of $\mathcal{L}_{0}$ is gapped.
    • $m=0$: As mentioned above, the presence of $E_{p}^{\alpha}(\alpha\neq 0)$ potentially rendering the relaxation rates gapless, contradicting the gapped relaxation rates for $\mathcal{L}_{0}$. To restore the physical observer argument, we propose the following conditions that could resolve the potential conflicts: the eigenspectrum of $\mathcal{L}_{\alpha\neq 0}$ is gapped; the eigenfunctions are localized within a finite $p$ region $p\leq p_{max}$ for $\mathcal{L}_{\alpha>0}$; the eigenfunctions are bounded in a region $p\geq p_{min}>0$ for $\mathcal{L}_{\alpha<0}$. As can be seen clearly, the argument imposes rather strong constraints, and furthermore, $\frac{1}{\gamma_{6}}\simeq\frac{p_{max}^{\alpha}}{\gamma^{\prime}_{6}}$ for $\alpha>0$ and $\frac{1}{\gamma_{6}}\simeq\frac{p_{min}^{\alpha}}{\gamma^{\prime}_{6}}$ for $\alpha<0$. We propose a testing method to verify or falsify the above assumed constraints: numerically solving the eigenspectrum of $\mathcal{L}_{\alpha\neq 0}$ for hard interactions [13], for instance, four fermions interaction in the low-energy electroweak sector.
    • $m\neq 0$: The conclusion for $\mathcal{L}_{\alpha>0}$ remains unchanged in accordance with $m=0$ case. For $\mathcal{L}_{\alpha<0}$, the nonzero mass regulates the relaxation timescales, ensuring that $\frac{E_{p}^{\alpha}}{\gamma^{\prime}_{6}}<\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$ as long as the eigenspectrum of $\mathcal{L}_{\alpha<0}$ is gapped. Hence, we arrive at the following statements: the eigenspectrum of $\mathcal{L}_{\alpha<0}$ is gapped, and $\frac{1}{\gamma_{6}}\simeq\frac{m^{\alpha}}{\gamma^{\prime}_{6}}$. Note that $m$ is the counterpart of $p_{min}$ in the aforementioned case.
@@ IV Pole or cut  — the non-analytical structures in retarded correlators
Two-point retarded correlation functions are crucial and insightful as they encapsulate rich information about the transport properties of many-body systems. Their analytical structures can reflect the characteristic properties of how thermal equilibrium is reached. For instance, the non-analytical structures — poles or cuts in Fourier space, governs the evolution behavior of the system: poles describe collective excitations evolving towards equilibrium corresponding to hydrodynamic modes, while the presence of cuts or non-hydrodynamic modes is closely linked to the emergence of hydrodynamic behavior and the applicability of hydrodynamics. Thus, the research into the analytical properties of retarded correlators is profound, which is initially explored by Romatschke in the weakly-coupled kinetic theory [7]. Romatschke’s findings highlight two key features in the analytical properties of the retarded correlators: the cuts are gapped corresponding to non-hydrodynamic modes, below which hydrodynamic poles dominate as long-lived degrees of freedom; the hydrodynamic poles cease to exist for some critical value of the wavenumber reminiscent of the phenomenon of onset transitions, which are successfully reproduced within the mutilated RTA model detailed in [34]. As a supplement, we note that the universal behavior of onset transitions has been reported for a long time in the context of the nonrelativistic kinetic theory using the mutilated model [19]. However, the AW RTA adopted in [7] is less well-founded than its non-relativistic counterpart, given that interactions tend to be softer in relativistic scenarios: the universal onset transitions in nonrelativistic systems may be rarely observed in relativistic systems due to varying degrees of interaction softness.
@@
Later on, Kurkela and Wiedemann reexamined the behavior of the retarded correlators, beginning with the parametrized energy-dependent RTA as described in Eq. (8), with a particular focus on the case when $\alpha=1$ [10]. Their conclusions, however, stand in contrast to those in [7]: there is no sharp onset of hydrodynamic behavior; the structure of cuts turns into the entire strip $\text{Im}\,\omega<0,-k\leq\text{Re}\,\omega\leq k$ from a gapped line given in [7]. They also state that the appearance of poles in the first (physical) Riemann sheet of retarded correlation functions is a matter of choosing a particular analytical continuation and thus cannot be related unambiguously to the onset of fluid dynamic behavior. However, as demonstrated in III.2, their model corresponding to Eq.(8) with $\alpha=1$ exhibits inconsistencies: it can only be seen as an incomplete truncation to its ultraviolet (UV) completion — the linearized Boltzmann equation, because an infinite number of slow modes are excluded. Although the analysis given in [10] is still illuminating, we choose to work with RTA only in the hard interaction case for theoretical consistence.
@@
As observed, the above two studies introduced above correspond to the models discussed in III.1 and III.2, thereby clarifying the relationship between Q1 and Q2. In this place, we want to give a general statement on the subject of pole/cut structure in the (stress-stress) retarded correlators. Before proceeding, let’s elaborate on how to derive the non-analytical structure if the interactions are soft. In this case, the eigenvalue spectrum of $\mathcal{L}_{0}$ is gapless and the typical stress-stress retarded correlation function is given by
[…]
@@
where we use $\gamma$ to denote the continuously distributed eigenvalues, and the weight function $\rho(\gamma)$ is nonzero in the integration range. This expression can be derived from the Fourier transform of the linearized Boltzmann equation Eq.(3) in the limit of vanishing $k$ [12, 33]. This expression clearly exhibits discontinuity as we shift from $\omega=-i\gamma+\epsilon$ to $\omega=-i\gamma-\epsilon$, indicating the presence of a branch-cut line extending from $-i\nu_{max}$ to 00.
@@
In the limit of vanishing mass and nonzero $k$, extracting the non-analytical structure becomes unmanageable. If $\cos\theta$, regarded as an operator in momentum-space functions, commute with $\mathcal{L}_{0}$, we can replace $\omega$ by $\omega-k\cos\theta$, where $\theta$ is the angle between spatial components of $p$ and $k$. Then we can cast $G_{R}(\omega)$ into
[…]
@@
[…]
@@
$\text{Im }\omega=-\gamma,\quad-k\leq\text{Re}\,\omega\leq k,\quad 0<\gamma<\infty,$ (31)
where the reason for $\nu_{max}=0$ is postponed to Eq.(32). Ultimately, we have successfully reproduce the branch-cut structure for nonzero $k$ given in [10] 222 [10] is based on the choice of $\mathcal{L}_{1}$, or more precisely, an incomplete truncation of $\mathcal{L}_{1}$, as shown in III.2. For completeness and consistency, a thorough study extending [35] to spatially inhomogeneous perturbations is warranted.. This accordance is expected because $[\cos\theta,\mathcal{L}_{\text{RTA}}]=0$.
However, a reminder should be given that Eq.(31) relies heavily on the commutation approximation for $\cos\theta$ and $\mathcal{L}_{0}$, which breaks down when $\mathcal{L}_{0}$ is the complete linearized Boltzmann collision operator rather than the RTA. In other words, the structure in Eq.(31) should be modified. For situations involving hard interactions, the extent of modification should be minimal as the RTA is well-defined in that case.
Based on the above statement, determining the non-analytical structures in the retarded correlation functions equates to solving for the eigenspectrum of $\mathcal{L}_{0}$, thereby interlinking Q1 and Q2. However, a contradictory question arises: both [12] and [13] demonstrate that the cut extends across the entire negative imaginary frequency axis, contrasting with the bounded region $[-i\,\nu_{max},0]$ inferred from mathematical analysis. The discrepancy vanishes because, in the context of a massless theory with soft interactions, the collision frequency becomes unbounded [32]
@@
Therefore, $\nu_{max}=\infty$ and these results are consistent with each other.
For clarity, we give a short summary of this section by reconsidering the interplay of nonzero particle mass $m$ and wavenumber $k$, and our statement can be summarized as follows:
@@
    • $m=0$: Romatschke’s analysis applies: the retarded correlators (including stress-stress correlator) exhibit gapped branch-cut lines
      $\text{Im }\omega=-\frac{1}{\tau_{R}},\quad-k\leq\text{Re}\,\omega\leq k$ (33)
      associated with nonhydrodynamic modes. Below this gap, hydrodynamic modes represent the sole low-energy degrees of freedom, indicating that hydrodynamics is well-defined in such scenarios.
    • $m\neq 0,\,k=0$: The branch-cut lines turn into branch points $\text{Im }\omega=-\frac{1}{\tau_{R}},\,\text{Re}\,\omega=0$.
    • $m\neq 0,\,k\neq 0$: The non-analytical structures of the retarded correlators become complicated due to the interplay between the nonzero particle mass $m$ and wavenumber $k$. Upon inspecting the derivation detailed in [7], we find that if the particles are massive, the free-streaming is proportional to $\bm{v}\cdot\nabla\sim\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ depending on $p$. This could introduce additional complex structures into correlators after performing momentum integral, beyond the poles or gapped cuts predicted by Romatschke. For now, we cannot exclude the possibility of closing the window where hydrodynamic modes are the only long-lived degrees of freedom. In the limit of vanishing mass or $k$, the conclusions align with the above two cases.
  • Soft interactions. The RTA does not constitute a well-defined model. The simplifying assumptions should at least retain the eigenvalues and eigenfunctions of $\mathcal{L}_{0}$ near the origin, specifically within the region $[0,\nu_{max}]$.
    • $m=0,\,k=0$: In this case, the stress-stress correlator possesses a branch cut described by
      $-\infty<\text{Im }\omega<0,\quad\text{Re}\,\omega=0,$ (34)
      which matches the conclusion given in [12]. Strictly speaking, the author focus on the discussions on the symmetrized 2-point function therein, but the retarded correlation function can be related to it through KMS relation. At weak coupling where the relevant frequencies will be suppressed by powers of the coupling, the result aligns with Eq.(34), which is further confirmed in a related study discussing $\mathcal{L}_{1}$ [35].
    • $m\neq 0,\,k=0$: The branch-cut of the stress-stress correlator turns into
@@
      where $\lim\limits_{m\rightarrow 0}\nu_{max}(m)\rightarrow\infty$. Therefore, the above result for $m=0,\,k=0$ can be recovered in the limit of vanishing $m$.
    • $k\neq 0$: The complication arising from $\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ also exists in this scenario, and the issue should be trickier. Even if $m=0$, it is still difficult to reach a concise-form conclusion due to the non-commutativity of $\cos\theta$ and $\mathcal{L}_{0}$.
Although failing to reach a definite conclusion in the cases involved with nonzero $k$ or/and $m$, we plan to numerically solve the issue following techniques developed in [13], which is left to a future work. For instance, for soft interactions with nonzero $k$ or/and $m$, we can examine the simplest interaction case, scalar $\phi^{4}$ theory with leading-order interaction. As a representative theory within the category of soft interactions, this calculation is expected to reveal the property of non-analytical structures qualitatively. Our numerical efforts following this direction are in progress.

### v3

@@ III.4 Hard interaction VS soft interaction
$\partial_{t}\tilde{\chi}(x,p)+\bm{v}\cdot\nabla\tilde{\chi}(x,p)=-\mathcal{L}_{0}[\tilde{\chi}]\equiv-v(p)\tilde{\chi}(x,p)+K[\tilde{\chi}$], (23)
@@
$K[\tilde{\chi}$] $\equiv f^{1/2}_{eq}(p)p_{0}^{-1}\int{\rm dP^{\prime}}{\rm dP_{1}}{\rm dP_{2}}f_{eq}(p^{\prime})W_{p,p^{\prime}\to p_{1},p_{2}}\big(\frac{\tilde{\chi}(x,p_{1})}{f^{1/2}_{eq}(p_{1})}+\frac{\tilde{\chi}(x,p_{2})}{f^{1/2}_{eq}(p_{2})}-\frac{\tilde{\chi}(x,p^{\prime})}{f^{1/2}_{eq}(p^{\prime})}\big),$ (25)
where $\nu(p)$ represents the collision frequency and we introduce a compact operator $K$. Here we maintain the same notations as in the previous context to avoid complicating the overall notation system, although the expansion Eq.(22) looks quite different. Furthermore, it is immediately observable that Eq.(23) shares the similar form as Eq.(7). As a result, they should also share the similar formal solution $\tilde{\chi}(t)=e^{-\mathcal{L}_{0}t}\tilde{\chi}(0)$ when the system is homogeneously perturbed. In both cases, whether considering Eq.(23) or Eq.(7), $\chi$ or $\tilde{\chi}$ should be regarded as the perturbation above the same equilibrium state. Given that both Eqs.(23) and (7) are derived from the same Boltzmann equation through linearization, we would not expect the eigenspectrum for $\mathcal{L}_{0}$ in Eq.(23) to differ from that in Eq.(7). If this were the case, the decay behavior of the perturbation would exhibit significant differences, signifying the physical inconsistency. This is another successful application of the observer argument from the previous subsection. Therefore, we argue that the eigenspectrum of $\mathcal{L}_{0}$ should be identical in Eqs.(23) and (7), and this is why we maintain the same notations. Then let’s come to the core of this subsection:
@@
  • Assume that $\exists\gamma>-2,\,0\leq\beta<\gamma+2,\,B>0$ and $c_{0}>0$ , so that $\sigma(g,\Theta)>B\frac{g^{\beta+1}}{c_{0}+g}\sin^{\gamma}\Theta$, then $\nu(p)>\nu_{0}(p_{0}/m)^{\beta/2}$ where $\nu_{0}$ is a constant, the interaction is hard.
@@
  • Assume that $\exists 0<\alpha<4,\,\gamma>-2$ and $B^{\prime}>0$, so that $\sigma(g,\Theta)<B^{\prime}g^{-\alpha}\sin^{\gamma}\Theta$, then $\nu(p)<\nu_{0}(p_{0}/m)^{-\epsilon/2}\leq\nu_{0}$, the interaction is soft, where
@@
    and $\nu_{0}$ is a constant,
where the proofs of these theorems can be found in [37].
@@
Analyzing the collision frequency $\nu(p)$ enables us to distinguish between relativistic soft and hard interactions. Mathematical analysis indicates that relativistic interactions tend to be softer than their nonrelativistic counterparts [36, 37]. It can be also proved that for soft interactions $\mathcal{L}_{0}$ is a bounded operator with the eigenspectrum $\nu_{max},0]$, whereas for hard interactions, $\mathcal{L}_{0}$ is unbounded, featuring an eigenspectrum of $\infty,-\nu_{min}]$, where $\nu_{max}$ and $\nu_{min}$ can be estimated by the extrema of $\nu(p)$. Given the eigenspectrum properties presented here, we conclude that the AW RTA is only well-justified in the case of hard interactions. For example, in the Weinberg-Salam theory at low energies [28], e.g., four fermions interaction in the electroweak sector well below the gauge boson masses, the typical differential cross section behaves as $\sigma(g,\Theta)\sim s=g^{2}+4m^{2}>\frac{g^{3}}{c_{0}+g}\geq\frac{g^{3}}{c_{0}+g}\sin\Theta$ with the conditions $\gamma=1,\beta=2,B=1$ in Theorem 1. Consequently, the model of a gas consisting of elastically colliding neutrinos in the low-energy limit permits an RTA description of transport phenomena.
As most relativistic interactions are soft according to Theorem 2, the AW RTA lacks a solid foundation in the majority of cases. For instance, in the extensively studied scalar $\phi^{4}$ theory, the interaction is soft as exhibited in the previous section. This softness results in a branch-cut structure that extends across the entire negative imaginary axis in the retarded correlators [13, 39].
There is one comment left. When interactions are too complicated so that their classification as hard or soft cannot be easily known, we recommend employing the finite-element method described in [14] for analyzing the eigenspectrum properties. By implementing the outlined procedures therein, we can establish the dictionary between various field theories and the eigenspectrum structure within their linearized kinetic description. In this sense, we can exhaust all commonly used interactions admitting an RTA description.
@@ IV Pole or cut  — the non-analytical structures in retarded correlators
Two-point retarded correlation functions are crucial and insightful as they encapsulate rich information about the transport properties of many-body systems. Their analytical structures can reflect the characteristic properties of how thermal equilibrium is reached. For instance, the non-analytical structures — poles or cuts in Fourier space, govern the evolution behavior of the system: poles describe collective excitations evolving towards equilibrium corresponding to hydrodynamic modes, while the presence of cuts or non-hydrodynamic modes is closely linked to the emergence of hydrodynamic behavior and the applicability of hydrodynamics. Thus, the research into the analytical properties of retarded correlators is profound, which is initially explored by Romatschke in the weakly-coupled kinetic theory [7]. Romatschke’s findings highlight two key features in the analytical properties of the retarded correlators: the cuts are gapped corresponding to non-hydrodynamic modes, below which hydrodynamic poles dominate as long-lived degrees of freedom; the hydrodynamic poles cease to exist for some critical value of the wavenumber reminiscent of the phenomenon of onset transitions, which are successfully reproduced within the mutilated RTA model detailed in [40]. As a supplement, we note that the universal behavior of onset transitions has been reported for a long time in the context of the nonrelativistic kinetic theory using the mutilated model [22]. However, the AW RTA adopted in [7] is less well-founded than its nonrelativistic counterpart, given that interactions tend to be softer in relativistic scenarios: the universal onset transitions in nonrelativistic systems may be rarely observed in relativistic systems due to varying degrees of interaction softness.
@@
Later on, Kurkela and Wiedemann reexamined the behavior of the retarded correlators, beginning with the parametrized energy-dependent RTA as described in Eq. (8), with a particular focus on the case when $\alpha=1$ [11]. Their conclusions, however, stand in contrast to those in [7]: there is no sharp onset of hydrodynamic behavior; the structure of cuts turns into the entire strip $\text{Im}\,\omega<0,-k\leq\text{Re}\,\omega\leq k$ from a gapped line given in [7]. They also state that the appearance of poles in the first (physical) Riemann sheet of retarded correlation functions is a matter of choosing a particular analytical continuation and thus cannot be related unambiguously to the onset of fluid dynamic behavior. However, as demonstrated in III.2, their model corresponding to Eq.(8) with $\alpha=1$ exhibits inconsistencies: it can only be seen as an incomplete truncation to its ultraviolet (UV) completion — the linearized Boltzmann equation, because an infinite number of slow modes are excluded. Although the analysis given in [11] is still illuminating, we choose to work with RTA only in the hard interaction case for theoretical consistency.
@@
As observed, these two studies introduced above correspond to the models discussed in III.1 and III.2. In this place, we want to give a general statement on the topic of non-analytical structures in the (stress-stress) retarded correlators. Before proceeding, let’s elaborate on how to derive the non-analytical structure if the interactions are soft. In this case, the eigenvalue spectrum of $\mathcal{L}_{0}$ is gapless and the typical stress-stress retarded correlation function is given by
$G_{R}(\omega)=\int^{\nu_{max}}_{0}d\gamma\frac{\rho(\gamma)}{\omega+i\gamma}$ (27)
@@
where we use $\gamma$ to denote the continuously distributed eigenvalues, and the weight function $\rho(\gamma)$ is nonzero in the integration range. This expression can be derived from the Fourier transform of the linearized Boltzmann equation Eq.(3) in the limit of vanishing $k$ [13, 39]. This expression clearly exhibits discontinuity as we shift from $\omega=-i\gamma+\epsilon$ to $\omega=-i\gamma-\epsilon$, indicating the presence of a branch-cut line extending from $i\nu_{max}$ to 0.
@@
In the limit of vanishing mass and nonzero $k$, extracting the non-analytical structure becomes unmanageable. We first assume that $\cos\theta$, regarded as an operator in momentum-space functions, commute with $\mathcal{L}_{0}$, then we can replace $\omega$ by $\omega-k\cos\theta$, where $\theta$ is the angle between spatial components of $p$ and $k$. Then we can cast $G_{R}(\omega)$ into
$G_{R}(\omega,k)\sim\int_{-1}^{1}d\cos\theta\int^{\nu_{max}}_{0}d\gamma\frac{\rho(\gamma)}{\omega-k\cos\theta+i\gamma},$ (28)
@@
$G_{R}(\omega,k)\sim\frac{1}{k}\int^{\nu_{max}}_{0}d\gamma\rho(\gamma)\big(\log(\omega+k+i\gamma)-\log(\omega-k+i\gamma)\,\big).$ (29)
@@
$\text{Im }\omega=-\gamma,\quad-k\leq\text{Re}\,\omega\leq k,\quad 0<\gamma<\nu_{max},$ (30)
where $\nu_{max}=\infty$ for a massless theory (see Eq.(31) ). Ultimately, the resulting branch-cut structure seemingly reproduces the result for nonzero $k$ given in [11] 222[11] is based on the choice of $\mathcal{L}_{1}$, or more precisely, an incomplete truncation of $\mathcal{L}_{1}$, as shown in III.2. For completeness and consistency, a thorough study extending [38] to spatially inhomogeneous perturbations is warranted..
However, a reminder should be given that Eq.(30) relies heavily on the commutation approximation for $\cos\theta$ and $\mathcal{L}_{0}$, which breaks down when $\mathcal{L}_{0}$ is the complete linearized Boltzmann collision operator. In other words, the structure in Eq.(30) should be modified. For situations involving hard interactions, the extent of modification should be minimal as the RTA can be seen as a good approximation to the complete linearized collision operator in that case.
Based on the above statement, determining the non-analytical structures in the retarded correlation functions equates to solving for the eigenspectrum of $\mathcal{L}_{0}$. However, a contradictory question arises: both [13] and [14] demonstrate that the cut extends across the entire negative imaginary frequency axis, contrasting with the bounded region $i\,\nu_{max},0]$ inferred from mathematical analysis. The discrepancy vanishes because, in the context of a massless theory with soft interactions, the collision frequency becomes unbounded [37]
@@
Therefore, $\nu_{max}=\infty$, and these results are consistent with each other.
For clarity, we give a summary of this section by reconsidering the interplay of nonzero particle mass $m$ and wavenumber $k$, and our statement can be summarized as follows:
@@
    • $m=0$ or $k=0$: Romatschke’s analysis applies: the retarded correlators (including stress-stress correlator) exhibit a gapped branch-cut line (only two endpoints are branch points)
      $\text{Im }\omega=-\frac{1}{\tau_{R}},\quad-k\leq\text{Re}\,\omega\leq k,$ (32)
      associated with nonhydrodynamic modes. Below this gap, there is a window where hydrodynamic modes become the long-lived degrees of freedom.
    • $m\neq 0,\,k\neq 0$: The non-analytical structures of the retarded correlators become complicated due to the interplay between the nonzero particle mass $m$ and wavenumber $k$. Upon inspecting the derivation detailed in [7], we find that if the particles are massive, the free-streaming term is proportional to $\bm{v}\cdot\nabla\sim\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ depending on $p$. This could introduce additional complex structures into correlators after performing momentum integral, beyond the poles or gapped cuts predicted by Romatschke. We leave this discussion to future work. For now, we can only conclude that the non-analytical structure remains in the form of Eq.(32). The remarkable distinction is that the points lying within the branch-cut line are all branch points.
  • Soft interactions. The RTA does not constitute a well-defined model. The simplifying assumptions should at least retain the eigenvalues and eigenfunctions of $\mathcal{L}_{0}$ near the origin, specifically within the region $0,\nu_{max}]$.
    • $k=0$: In this case, the stress-stress correlator possesses a branch cut described by
@@
      which in the limit of vanishing mass matches the conclusion drawn in [13]. Strictly speaking, the author focuses on the discussions on the symmetrized 2-point function therein, but the retarded correlation function can be related to it through KMS relation. At weak coupling where the relevant frequencies will be suppressed by powers of the coupling, the result aligns with Eq.(33), which is further confirmed in a related study discussing $\mathcal{L}_{1}$ [38].
    • $k\neq 0$: The complication arising from $\frac{pk\cos\theta}{\sqrt{p^{2}+m^{2}}}$ also exists in this scenario, and the issue should be trickier than the case of massive RTA. Even if $m=0$, it is still difficult to reach a concise-form conclusion due to the non-commutativity of $\cos\theta$ and $\mathcal{L}_{0}$. Therefore, the non-analytical structures don’t exhibit regular shapes like Eq.(30).
In the end of this section, we would like to add several comments as follows:
— Although failing to reach a definite conclusion in the cases involved with nonzero $k$ or/and $m$, we plan to numerically solve the issue following finite-element analysis given in [14], which is left to future work. For instance, for soft interactions, we can examine the simplest interaction case, scalar $\phi^{4}$ theory with leading-order interaction, which is expected to reveal the property of non-analytical structures qualitatively.
— If Eq.(30) holds, hydrodynamic modes are completely embedded in this strip, leading to the conclusion that nonhydrodynamic modes related to free-streaming dynamics are long-lived degrees of freedom compared to hydrodynamic modes. According to our proposed non-commutativity between $\cos\theta$ and $\mathcal{L}_{0}$, Eq.(30) would be modified so that there could be a window where hydrodynamic modes dominate as long-lived modes. In this sense, the physical picture may change by reconsidering the impact of the non-commutativity.
— In this script (likewise in [27]), we focus on the gapped/gapless property of nonhydrodynamic modes. By using the terminology “dominant” or “dominate”, we refer to the comparison of their lifetime. In practical cases, the dominant role should be assessed by combining their corresponding residues or discontinuity, which encode their contribution to the retarded correlation functions. This issue can be addressed using numerical calculations outlined in the first comment.